HOCKEY teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 57 points. They won 12 more games than they tied. How many wins and ties did the team have?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

HOCKEY teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 57 points. They won 12 more games than they tied. How many wins and ties did the team have?
------------------------------------------------------------
Quantity Eq.: w = t + 12
Points Eq.: 2w + 1t = 57
---------------------------------
Substitute for "w" and solve for "t":
2(t+12) + t = 57
2t + 24 + t = 57
3t = 33
t = 11 (# of games the team tied)
---
Since w = t + 12, w = 11 + 12 = 23 (# of games the team won)
==========================
Cheers,
Stan H.

Answer by richwmiller(17219)

w=12+t
2w+t=57
2(12+t)+t=57
24+2t+t=57
3t=33
t=11
w=t+12
w=23
check
2*23+11=57
46+11=57